### v1

# MTQA: Matrix of Thought for Enhanced Reasoning in Complex Question Answering

Abstract
Complex Question Answering (QA) is a fundamental and challenging task in NLP. While large language models (LLMs) exhibit impressive performance in QA, they suffer from significant performance degradation when facing complex and abstract QA tasks due to insufficient reasoning capabilities. Works such as Chain-of-Thought (CoT) and Tree-of-Thought (ToT) aim to enhance LLMs’ reasoning abilities, but they face issues such as in-layer redundancy in tree structures and single paths in chain structures. Although some studies utilize Retrieval-Augmented Generation (RAG) methods to assist LLMs in reasoning, the challenge of effectively utilizing large amounts of information involving multiple entities and hops remains critical. To address this, we propose the Matrix of Thought (MoT), a novel and efficient LLM thought structure. MoT explores the problem in both horizontal and vertical dimensions through the ”column-cell communication” mechanism, enabling LLMs to actively engage in multi-strategy and deep-level thinking, reducing redundancy within the column cells and enhancing reasoning capabilities. Furthermore, we develop a fact-correction mechanism by constructing knowledge units from retrieved knowledge graph triples and raw text to enhance the initial knowledge for LLM reasoning and correct erroneous answers. This leads to the development of an efficient and accurate QA framework (MTQA). Experimental results show that our framework outperforms state-of-the-art methods on four widely-used datasets in terms of F1 and EM scores, with reasoning time only 14.4% of the baseline methods, demonstrating both its efficiency and accuracy. The code for this framework is available at https://github.com/lyfiter/mtqa.
Introduction
The human transition from learning to practice generally entails comprehending and reasoning over diverse information sources, followed by making decisions that align with—or even extend beyond—the original information. Recent studies have demonstrated that large language models (LLMs) can replicate this learning-to-practice paradigm in complex domains such as intelligence analysis (Wu et al. 2024) and creative writing (Qin et al. 2024). Complex question answering (QA) is a foundational and critical downstream task for LLMs (Zhuang et al. 2023); it demands that a QA system command extensive knowledge and perform sophisticated reasoning over that knowledge to deliver satisfactory answers.
Figure 1: Comparison of Chain of Thought, Thought of Tree, and MoT for Complex Question Answering.
Recently, large language models (LLMs) have become one of the most promising and efficient solutions in the QA field, thanks to their massive pre-trained knowledge and the reasoning abilities they autonomously generate during training(Guo et al. 2025). However, some studies indicate that LLMs exhibit significant shortcomings in complex question-answering reasoning tasks. Various research efforts have attempted to enhance LLMs’ reasoning performance using thought structures. For example, the Chain-of-Thought (CoT) framework(Wei et al. 2022) breaks down the reasoning process of LLMs into a series of ordered steps, thereby improving their reasoning capabilities. However, CoT suffers from a single reasoning path, limiting its ability to explore multiple reasoning paths within the solution space, and may also result in incorrect ideas that cannot be corrected in time, leading to deviations. Tree-of-Thought (ToT)(Yao et al. 2023a) organizes the reasoning process into a hierarchical tree structure, allowing the model to reason across multiple layers and branches. However, its node selection mechanism is inefficient and often produces redundant nodes, leading to resource wastage and ineffective reasoning.
Furthermore, some studies have employed Retrieval-Augmented Generation (RAG) to assist LLMs in answering questions, enhancing them with external knowledge through retrieval and generation, leading to more accurate and reliable answers(Cuconasu et al. 2024; Fan et al. 2024). Currently, mainstream reasoning methods incorporate RAG into thought structures to provide LLMs with the ability to validate facts. RAT(Wang et al. 2024) refines the step-by-step reasoning process of CoT by adding a RAG correction at each reasoning stage to ensure the reasoning trajectory remains on track. RATT(Zhang et al. 2025) combines the multi-branch, multi-level reasoning of tree structures with RAG’s external knowledge to correct factual errors, thereby strengthening LLMs’ reasoning capabilities. However, these works use vanilla RAG as the fact-validation module, which limits its ability to assist in reasoning and correct facts. In complex question-answering tasks, especially in open-ended tasks involving multiple entities and multi-hop information, the retrieved knowledge often contains irrelevant or even erroneous information, which misguides the LLMs’ reasoning process. The quality of the factual verification knowledge provided by RAG becomes one of the key factors in guiding LLMs’ reasoning.
To address the aforementioned issues and improve the reliability and accuracy of LLMs in complex question answering tasks, we aim to develop a method that efficiently utilizes existing knowledge and performs multi-branch reasoning with low redundancy to generate accurate answers. In terms of efficiently utilizing existing knowledge, we focus on exploring the representation of knowledge retrieved by RAG to assist LLMs in answering questions. In the context of multi-branch and low-redundancy reasoning, we aim for the model to autonomously adopt different perspectives to understand the existing knowledge and perform reasoning that leads to an answer close to the correct one, all while using the same initial knowledge.
To achieve these goals, we propose a novel Matrix of Thought (MoT) approach to ensure that the reasoning process for complex question-answering tasks is multi-branching and low-redundancy. As shown in Figure 1, in MoT, our RAG retrieves entity-relation-entity triples that match the question, along with their corresponding original text segments, which are then constructed into knowledge units to serve as auxiliary knowledge for initiating reasoning and fact verification. Additionally, we introduce a column cell communication mechanism that actively helps the LLM explore alternative perspectives for solving the problem, thereby reducing the redundancy in the generated reasoning process within the column. Then, MoT combines the various branching strategies with RAG’s auxiliary knowledge for summarization and optimization, using this as the reasoning initiation knowledge for the next column cell, thereby approaching the most promising answer.
The contributions of this paper are summarized as follows:
1. We propose an innovative large model reasoning paradigm, MoT, which addresses the issues of information redundancy within the same layer of tree structures and the single decision path in chain structures. MoT enhances the flexibility and adaptability of reasoning at a minimal cost.
2. We have developed a novel framework for complex QA tasks, the Matrix of Thought Question Answering (MTQA) framework. This framework innovatively combines KG triples with their corresponding original text into knowledge units for knowledge enhancement and correction. By integrating the MoT reasoning paradigm, it achieves a fully automated QA pipeline, enabling efficient reasoning on the problem from both breadth and depth.
3. We conducted a series of experiments to validate the efficiency and reliability of our method in complex question answering across various domains. The results demonstrate that our method holds a clear advantage over existing approaches.
Related Work
Thought Structures for LLMs
The thought structures proposed (Ding et al. 2023; Minaee et al. 2024) is a mainstream prompt engineering approach that guides LLMs to engage in more logical thought processes, thereby improving the quality of generated responses. Chain of Thought (Wei et al. 2022) is a representative work that guides LLMs to generate intermediate reasoning steps within the prompt, thereby steering the large model through the reasoning process. Wang et al. developed the Self-consistency with Chain of Thought (CoT-SC) method (Wang et al. 2023), which extends CoT by solving a problem using multiple independent and distinct chains of thought. The most reliable answer is then selected as the final output response. The Tree of Thoughts (ToT) (Yao et al. 2023a) abstracts the reasoning process of a problem into a tree search. By constructing multiple branches and evaluating their contributions to solving the problem, ToT guides the LLM in reasoning, enabling a broad and global exploration of the solution space.
Current mainstream research focuses on CoT and ToT; however, the inherent properties of their structures present performance bottlenecks, leading to issues such as single reasoning paths, generation of redundant information, and the risk of deviating from the reasoning trajectory due to incorrect ideas during inference. One of the key focuses of our research is to develop a more efficient and reliable thought structure paradigm, enabling LLMs to perform exceptionally well in QA tasks and beyond within this new framework.
Retrieval-Augmented Generation (RAG) for LLM on QA
When LLMs solve extremely complex tasks, they suffer from ”complete accuracy collapse”(Shojaee*† et al. 2025), significant performance degradation, and LLMs hallucination problems due to insufficient reasoning capabilities(Sriramanan et al. 2024). Retrieval-Augmented Generation (RAG)(Lewis et al. 2020) is a pivotal technique for mitigating such hallucinations and enhancing the quality of LLM outputs. Retrieval-Augmented Generation (RAG) stores task-relevant knowledge in an external repository and, when the LLM is invoked, retrieves pertinent information to serve as supplementary context. GraphRAG (Edge et al. 2024) extends this paradigm by incorporating knowledge graphs (KGs): an LLM produces concise community-level summaries that both abstract the input documents and provide graph indices; each community then generates its response independently, and the individual outputs are subsequently aggregated into a global answer. Recent studies have begun to integrate explicit reasoning paradigms into the RAG workflow. Retrieval-Augmented Thought (RAT)(Wang et al. 2024) merges the RAG retrieval step directly into the chain-of-thought process, allowing the model to iteratively refine and correct its reasoning step by step. Retrieval-Augmented Thought Tree (RATT) (Zhang et al. 2025) adopts a tree-structured reasoning framework in which each node invokes RAG for factual verification, thereby enhancing the logical soundness and factual reliability of LLM reasoning.
However, current research often treats the retrieved knowledge as various fragmented KG triples or text, which do not align well with the preferences of LLMs. Additionally, there has been limited work on designing RAG systems specifically for assisting thought structure reasoning. This paper focuses on studying the optimal representation of the knowledge retrieved by RAG to aid in reasoning within thought structures, enabling LLMs to efficiently utilize the knowledge for reasoning and derive the best possible answers.
Methodology
This section presents the Matrix of thought Question Answering (MTQA) framework in detail. As illustrated in Figure 2, MTQA is constructed atop the novel MoT reasoning paradigm to tackle complex QA tasks. The framework is designed to make fuller use of existing knowledge and to equip LLMs with both breadth and depth when exploring the problem’s solution space. Our method improves the reliability and comprehensiveness of LLM-generated answers from two aspects: adopting a superior knowledge representation to enhance the LLM’s comprehension. And employing a powerful thought structures to strengthen the LLM’s reasoning capability. The framework consists of two parts: (1) retrieval-augmentation, (2) Matrix of thought.
Figure 2: Framework of MTQA. Given an input including a query and source documents, the LLM retrieves relevant knowledge units from both knowledge graphs and text databases. These knowledge units are then integrated into the Matrix of Thought (MoT), where multiple reasoning strategies are explored in parallel through column-cell communication. The reasoning process is enhanced by retrieval-augmented knowledge, and the final answer is refined through correction and summarization.
Problem Definition
A question-answering task comprises a given question $Q$ and a set of input documents $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{n}\}$, where $n$ denotes the number of input documents. Both the question and the input documents are expressed in natural language. $D$ is the auxiliary factual knowledge that supports the large language model $L$ in answering $Q$ and thus obtaining the answer $A$. This factual knowledge can take the form of textual passages, triples in a knowledge graph, or Internet web page URLs. This paper focuses on treating text paragraphs as auxiliary factual knowledge. Using the retriever $R$, we retrieve from $D$ the factual knowledge set $\mathcal{K}=\{k_{1},k_{2},\ldots,k_{m}\}$ that matches the current input $X$ to be verified by $L$, where $m$ represents the top-$m$ text blocks ranked by their degree of match with $X$.
In MTQA, our objective is to construct a matrix of thought $M$ that can carry out the widest and deepest possible inference within the solution space of question $Q$, thereby identifying the optimal answer $A$.
Retrieval-Augmented
Step 1: Knowledge Database construction.In MTQA, the first step is to build a retrievable external knowledge base and partition the given set of documents into chunks whose size is suitable for processing by the LLM:
$C=\{c_{1},\,c_{2},\,c_{3},\,\dots,\,c_{k}\},\quad k=\left\lceil\frac{N_{D}}{C_{\text{len}}}\right\rceil$ (1)
Where $C$ denotes the set of blocks after the document group $D$ is chunked, $k$ represents the number of blocks, $N_{D}$ is the total number of words in the document group $D$, and $C_{len}$ is the length of the set text block.
Subsequently, we use LLM to recognize and extract entities (e.g., dates, places, people, events, etc.) and the relationships between them according to set cues (Detailed prompt are presented in the Appendix), generating a knowledge graph $G$ that comprehensively encompasses the content of the document, which consists of the extracted entities $v$ with their relationships with each other $\varepsilon$, and reacts to the connections and focuses in the whole document set.
First, llm embeds the chunked text block $C_{i}$ after filling it to the prompt template:
$L_{c_{i}}=\operatorname{embed}\!\bigl{(}\operatorname{template}(C_{i},x)\bigr{)}$ (2)
Where $x$ represents the template for performing entity and relationship extraction, $C_{i}$ represents the text block, and $L_{c_{i}}$ represents the generated embedding vectors subsequently input to the LLM for comprehension. The LLM performs recognition and extraction operations to extract the entity $v$ and the relation $\varepsilon$ from $L_{c_{i}}$.
$v,\;\varepsilon=\text{Extraction \& identification}\!\left(\sum_{i=1}^{k}L_{c_{i}}\right)$ (3)
Subsequently, we develop guidance llm to construct a key-value pair $(k,v)$ for each recognized and extracted entity $v$ and the relation $\varepsilon$ between them. The keys are composed of individual words or short phrases and are later used as query terms to retrieve the corresponding entities and relations. For entities, their names function as keys; for relations, the LLM summarizes the text that describes the link between two entities to generate multiple keys. The associated values are the source passages in the external knowledge base from which those entities and relations are drawn. The key–value mechanism both accelerates retrieval and preserves direct access to the source passages in the external knowledge base. During the generation of entities and their corresponding relations, the LLM augments each item with an explanatory description, allowing subsequent LLM-based reasoning modules to exploit this enriched knowledge more effectively for refinement and enhancement.
$\hat{v},\;\hat{\varepsilon}=\operatorname{Deduplication}\!\bigl{(}v,\;\varepsilon\bigr{)}$ (4)
After completing the extraction of entity $v$ and relation $\varepsilon$, we perform a de-redundancy operation to identify and merge the duplicates of $v$ and $\varepsilon$ to obtain the unique $\hat{v}$ and $\hat{\varepsilon}$ in the whole knowledge graph, which improves the overhead of the subsequent operations on the knowledge graph and ensures that there is no redundancy in the retrieved knowledge.
$G=\bigl{(}\hat{v},\;\hat{\varepsilon}\bigr{)}$ (5)
[…]
After constructing the knowledge graph, we consider the incremental update operations for the knowledge graph when the document group is updated. For the newly added document group $D^{\prime}$, we follow the aforementioned steps to extract entities and relations, obtaining a new subgraph $G^{\prime}=(V^{\prime},\varepsilon^{\prime})$. Next, the entities and relations in $G$ and $G^{\prime}$are merged by taking their union, thereby combining $G^{\prime}$ with $G$.
The key-value structure of the knowledge graph that we construct as the external knowledge base for MTQA offers two advantages:
(1) Compared to indexing the original text passages, the knowledge graph stores structural information, providing the MTQA framework with multi-hop reasoning capabilities to integrate global information, enabling LLMs to understand knowledge across text blocks.
(2) The key-value structure strongly associates the knowledge graph with its source texts, facilitating efficient retrieval while enabling the construction of knowledge units that combine knowledge graph triples with source text-enhanced knowledge.
Step 2: Retrieval knowledge through queries. After constructing the external knowledge base, we build the retrieval function mechanism to match the knowledge that best supports the query. We use a dual retrieval mechanism (Edge et al. 2024) to extract relevant information from the numerous and complex intertwined dependencies of entities in the external knowledge base. This includes detailed queries and summary queries. Detailed queries focus on specific details and require precise identification of relevant entities or relations for factual judgment or providing key evidence. Summary queries, on the other hand, are broader in scope, summarizing topics or abstracts. They typically involve searching across multiple entity-relation triples or even text blocks, and are used to answer open-ended or comprehensive questions. We perform retrieval by combining graph structure information with vector representations. The retrieval process is mainly divided into three steps:
(1) Extract the keywords from the query $Q$, distinguishing between detailed keywords $k^{(d)}$ and generalized keywords $k^{(g)}$.
(2) Match the extracted detailed keywords with the entity keys in the vector database, and match the generalized keywords with the relation keys.
(3) Further retrieve the neighboring nodes (one-hop and two-hop adjacent nodes) in the knowledge graph for the matched entities and relations.
The retrieved knowledge consists of entities (nodes of the knowledge graph), relations (edges of the knowledge graph), and the original text (indexed by the key-value pairs of nodes and edges), forming knowledge units, as shown in Figure 2. The knowledge units simultaneously contain structured information, supporting multi-hop reasoning and emphasizing key information through knowledge graph triples(Saxena, Tripathi, and Talukdar 2020), as well as undistorted, broader content original (Lewis et al. 2020) (such as complex metaphors, emotional expressions, etc.), which better aligns with the preferences of LLMs (Dong et al. 2025) in original text. Experimental results demonstrate that these knowledge units significantly enhance the reasoning performance of LLMs.
Matrix of Thought
Step 3: Reasoning enhance. Next, we start building the column units in MoT. A given query $Q$ is fed into the retrieval enhancement module to get the knowledge unit $KU_{1}$ that underpins the reasoning of LLMs:
$KU_{1}=\text{Retrieval-Augmented}(S(Q))$ (6)
Next, we fill the knowledge unit $KU_{1}$ and the query $Q$ into the prompt template, guiding the LLMs to leverage the information contained in the knowledge unit to enhance their understanding of the question. A strategy from shallow to deep is employed to guide the LLMs in generating the initial thought nodes:
$Tn_{1}^{1}=\text{LLM}(\text{template}(Q,KU_{1}))$ (7)
The initial nodes contain relatively shallow and indicative views and solutions to the question. MTQA will iteratively explore and refine these nodes in terms of breadth and depth, gradually evolving them so that the generated thought nodes converge towards the target of the question.
Step 4: Column cell communication mechanism. To more efficiently explore the breadth of the problem and implement a multi-branching strategy, we have developed the column cell communication mechanism. Specifically, in MoT, a portion of the paragraph that represents the strategy and direction of the previous thought node in the same column is extracted (typically the last few paragraphs of the thought node). This portion is then used as an example to guide the LLMs in generating the next thought node, encouraging a completely different line of thought and strategy.
$Tn_{m}^{n}=LLM\left(\mathrm{template}\left(Q,KU_{n},\alpha Tn_{m-1}^{n}\right)\right)$ (8)
Where $\alpha$ is the communication weight with range $0\leq\alpha\leq 1$. It denotes the proportion (rounded up) of paragraphs taken from node $Tn_{m-1}^{\,n}$. A value of 0 means there is no communication between the two thought nodes, and a value of $1$ means the entire content of $Tn_{m-1}^{\,n}$ is passed to $Tn_{m}^{\,n}$. This parameter determines the transparency of prior ideas to the LLMs. In MoT, the column-cell communication mechanism is essential and distinguishes MoT from ToT and related reasoning structures. Throughout MoT, we use a matrix $A$ to represent the communication weights. Given an $m\times n$ thought matrix, the weight matrix has dimension $(m-1)\times(n-1)$. The optimal value of this parameter is detailed in the Experiments section.
$A=\begin{bmatrix}\alpha_{m-1}^{\,1}&\cdots&\alpha_{m-1}^{\,n-1}\\ \vdots&\ddots&\vdots\\ \alpha_{1}^{\,1}&\cdots&\alpha_{1}^{\,n-1}\end{bmatrix}$ (9)
[…]
$KU_{n}^{\prime}=\text{Retrieval-Augmented}\!\left(S\!\left(Q,\sum_{i=1}^{m}Tn_{i}^{\,n}\right)\right)$ (10)
$S_{n}^{\,n}=\text{LLM}\!\bigl{(}\text{template}(Q,KU_{n}^{\prime})\bigr{)}$ (11)
The summary node corrects any fact-inconsistent strategies that appear among the thought nodes in the column, while preserving as much strategy diversity as possible. When generating the next column cell, the process builds on the current column’s summary node to further develop promising strategies. After a specified number of iterations, the final summary node is taken as the output of MTQA.
Algorithm 1 MTQA Framework
0: Question $Q$, Document set $\mathcal{D}=\{d_{1},d_{2},\dots,d_{n}\}$, Retriever $R$, LLM $L$, Knowledge graph $G$
0: Final Answer AA
1:Step 1: Knowledge Database Construction
[…]
3:for each chunk $c_{i}\in C$ do
4:  $L_{c_{i}}\leftarrow\operatorname{embed}(\operatorname{template}(c_{i},x))$
5:  $v,\varepsilon\leftarrow\text{Extraction \& Identification}(L_{c_{i}})$
6:end for
[…]
14:for each column $n$, row $m$ do
15:  $Tn_{m}^{n}\leftarrow L(\operatorname{template}(Q,KU_{n},\alpha Tn_{m-1}^{n}))$
16:  $A\in\mathbb{R}^{(m-1)\times(n-1)}$
17:end forStep 5: Refinement and Summarization
[…]
Discussion
In this chapter, we further elaborate on the objectives introduced in the Introduction. While controlling the framework’s runtime overhead, our method integrates external knowledge and refines it into knowledge units to help LLMs efficiently enhance reasoning and correct strategies. Moreover, by adopting a matrix structure and a column-cell communication mechanism, it encourages LLMs to proactively explore multiple dimensions of the solution space. The MTQA framework for complex question answering is an instance of the MoT paradigm. It is worth noting that the MoT paradigm abstracts the LLM reasoning process as a matrix. When the number of columns or rows of this matrix equals 1, MoT degenerates into a CoT structure with RAG-based correction. When the column-cell communication weight is 0, MoT degenerates into a ToT-like structure. This indicates that CoT and ToT are special cases of MoT, which confers greater flexibility and adaptability on MoT.
Experiment
Experimental Setup
In this chapter, we design two experiments to validate the effectiveness of MTQA.
(1) Multi-metric performance evaluation: we conduct comparative experiments between MTQA and existing advanced methods using two metrics, the F1 score and EM.
(2) Multidimensional win-rate evaluation: we prompt a powerful LLM to perform pairwise, multi-dimensional assessments between MTQA and each baseline model’s responses, generate an evaluation report, and select the superior answer to compute the win rate.
To implement the MTQA framework, we default to GPT-4o-mini as the LLM $L$. To ensure consistency with prior work (Trivedi et al. 2022), the dataset chunk size is set to 1200. We use the nano vector database for vector database operations. The hyperparameter settings for MoT are detailed in the Numerical Analysis subsection of this chapter. All experiments are conducted on a Linux server equipped with six NVIDIA A6000 GPUs.
Multi-metric performance evaluation
In this experiment, we use NaturalQuestions (Kwiatkowski et al. 2019) for single-hop QA evaluation, and HotpotQA (Yang et al. 2018) together with 2WikiMultihopQA (2WMQA) (Ho et al. 2020) for multi-hop QA evaluation. To ensure consistency with prior work, we sample 500 instances from each dataset. We use the F1 score and Exact Match (EM) as evaluation metrics for all three datasets. We compare MTQA with GPT-4o-mini under direct prompting, Zero-shot CoT(Kojima et al. 2022), Few-shot CoT(Wei et al. 2022), and ToT, and against RAG-based baselines including vanila RAG(Lewis et al. 2020), ReAct(Yao et al. 2023b), IRCoT(Trivedi et al. 2022), FLARE(Jiang et al. 2023), Self-Rag(Asai et al. 2024), SearChain(Xu et al. 2024), Rowen(Ding et al. 2024), SlimPLM(Tan et al. 2024), LightRAG(Edge et al. 2024), CoA(Zhang et al. 2024), RAT(Wang et al. 2024), and RATT(Zhang et al. 2025). For RATT, we configure its internal RAG to be identical to that of MTQA (the original work uses online web pages as the external knowledge base, which we replace with the knowledge units described above).
Dataset Metric NQ HotpotQA 2WMQA
F1 EM F1 EM F1 EM
Vanilla GPT 4o-mini 0.425 0.298 0.380 0.265 0.316 0.229
Zero-shot CoT 0.457 0.295 0.355 0.262 0.324 0.216
Few-shot CoT 0.447 0.298 0.376 0.255 0.363 0.227
ToT 0.456 0.305 0.368 0.267 0.345 0.237
Vanilla RAG 0.385 0.258 0.387 0.254 0.314 0.244
ReAct 0.335 0.212 0.390 0.270 0.305 0.204
IRCoT 0.344 0.216 0.361 0.232 0.318 0.202
FLARE 0.455 0.318 0.391 0.268 0.364 0.246
Self-Rag 0.387 0.270 0.357 0.220 0.311 0.210
SearChain 0.337 0.214 0.349 0.216 0.313 0.222
Rowen 0.452 0.286 0.382 0.240 0.307 0.212
SlimPLM 0.442 0.280 0.393 0.266 0.368 0.242
LightRAG 0.439 0.312 0.362 0.282 0.356 0.276
CoA 0.452 0.326 0.387 0.272 0.362 0.243
RAT 0.462 0.338 0.392 0.275 0.376 0.239
RATT 0.479 0.323 0.410 0.279 0.406 0.284
MTQA    0.510 $\triangle$3.1%    0.368 $\triangle$3.0%    0.452 $\triangle$4.2%    0.318 $\triangle$3.6%    0.454 $\triangle$4.8%    0.326 $\triangle$4.2%
Table 1: Overall results on three datasets for multi-metric performance evaluation. The existing best results are underlined and the best results are in bold. $\triangle$ indicates the improvement of our results over the state-of-the-art results.
Results Analysis: We compare MTQA with the above baselines, as shown in Table 1. Based on the evaluation metrics, we draw the following conclusions: First, MTQA outperforms all baselines. Compared with non-RAG baselines, our method increases the F1 score by 5.3%–9.1% and EM by 6.2%–8.9%. Relative to RAG-based and reasoning-structure methods, we still maintain at least a 3% lead in both F1 and EM, which is especially pronounced on multi-hop QA datasets such as HotpotQA and 2WMQA. These results demonstrate the effectiveness of the proposed MoT paradigm and knowledge units. Second, we observe that several RAG methods do not consistently surpass the non-RAG baselines. This may be due to inefficient retrieval mechanisms that introduce noise or erroneous evidence, thereby misleading the LLM’s responses. Finally, our method achieves substantial gains in both F1 and EM, indicating that the column-cell communication mechanism effectively mines multi-branch lines of reasoning and increases response breadth, thereby improving answer quality.
In the above experimental setup, we use the F1 score and EM as evaluation metrics. However, the gold answers in these datasets are typically short, while LLMs tend to produce logically structured paragraph-length responses (Li et al. 2022), and answers may be semantically similar yet lexically non-identical (Zhou et al. 2023). These factors render F1 and EM insufficiently rigorous and prevent them from evaluating the quality of non-standard answers in open-domain QA (e.g., describing the personality traits of a novel’s protagonist).
Multidimensional win-rate evaluation
To address the issues in Multi-metric performance evaluation, we build on the method of Guo et al. (Edge et al. 2024) and design a simple yet effective protocol to evaluate, in a multidimensional manner, the response quality of QA models for open-domain QA tasks. We select four datasets from UltraDomain (Qian et al. 2025), which is compiled from 428 English-language university textbooks and spans 18 disciplines. Specifically, we use the Biology, Law, Computer Science (CS), and Physics subsets, each containing millions of tokens.
Implementation details: To evaluate MTQA’s effectiveness on high-level sensemaking QA tasks, we adopt the generation procedure summarized by Edge et al(Edge et al. 2024), instructing the LLM to generate five users in the target domain, with each user providing five tasks. Each user is assigned domain-specific knowledge and attributes to elicit greater nuance in task and question formulation. For each task, the LLM generates ten questions that require understanding of the entire corpus. In total, each dataset yields 250 domain-aligned questions. For the responses produced by MTQA and all baselines, we instruct GPT-4o to conduct pairwise evaluations (MTQA vs. each baseline) along four dimensions, including:
Comprehensiveness: How much detail the answer provides to cover all aspects and details of the question.
Accuracy: How well the answer ensures factual correctness, data reliability, and logical consistency across different perspectives and insights.
Empowerment: How well the answer helps the reader understand the topic and make informed judgments.
Overall: An aggregate assessment of the preceding three indicators to determine the best overall answer.
Regarding the baselines, we select the graph-based RAG system LightRAG, as well as CoT-style methods with RAG correction (CoA and RAT) and a ToT-style method (RATT), to compute win rates against MTQA. In RAT and RATT, the RAG component is configured identically to that of MTQA. An example of the evaluation process is shown in Figure 10 in the appendix.
Metric Biology CS Legal Physics
LightRAG MTQA LightRAG MTQA LightRAG MTQA LightRAG MTQA
Comprehensiveness 5.6% 94.4% 4.8% 95.2% 8.8% 91.2% 5.2% 94.8%
Accuracy 6.4% 93.6% 4.0% 96.0% 4.8% 95.2% 5.6% 94.4%
Empowerment 3.6% 96.4% 6.8% 93.2% 6.0% 94.0% 7.6% 92.4%
Overall 4.4% 95.6% 3.6% 96.4% 5.2% 94.8% 5.2% 94.8%
CoA MTQA CoA MTQA CoA MTQA CoA MTQA
Comprehensiveness 13.6% 86.4% 9.6% 90.6% 11.2% 88.8% 17.2% 80.8%
Accuracy 15.6% 84.4% 13.6% 86.4% 13.6% 86.4% 12.0% 88.0%
Empowerment 17.2% 82.8% 15.6% 84.8% 16.4% 83.6% 18.8% 81.2%
Overall 15.2% 84.8% 10.4% 89.6% 12.4% 87.6% 13.2% 86.8%
RAT MTQA RAT MTQA RAT MTQA RAT MTQA
Comprehensiveness 17.2% 82.8% 10.8% 89.2% 16.4% 83.6% 18.8% 81.2%
Accuracy 17.6% 82.4% 25.2% 74.8% 23.6% 76.4% 22.4% 77.6%
Empowerment 19.6% 80.4% 22.8% 77.2% 16.4% 83.6% 20.8% 79.2%
Overall 16.4% 83.6% 11.2% 88.8% 15.2% 84.8% 18.4% 81.6%
RATT MTQA RATT MTQA RATT MTQA RATT MTQA
Comprehensiveness 16.8% 83.2% 12.0% 88.0% 22.4% 77.6% 29.6% 70.4%
Accuracy 34.4% 65.6% 32.0% 68.0% 16.4% 83.6% 32.0% 68.0%
Empowerment 24.0% 76.0% 19.2% 80.8% 24.8% 75.2% 19.2% 80.8%
Overall 18.4% 81.6% 15.2% 84.8% 18.4% 81.6% 15.2% 84.8%
Table 2: The winning rate of MTQA and its baseline in four datasets and four evaluation dimensions in the Multidimensional win-rate evaluation. The one with the higher winning rate is marked in red.
Results Analysis: As shown in Table 2, we compare the win rates between MTQA and the aforementioned baselines pairwise. It is evident that our method achieves a win rate of over 80% across most dimensions and in the overall assessment, further validating the effectiveness of the MoT approach. Among them, the non-reasoning-structure method (LightRAG) achieves a win rate of no more than 10% across the four datasets. This further supports our viewpoint that, under similar knowledge enhancement, reasoning structures play a critical role in the utilization of knowledge and the logical reasoning ability, significantly influencing the overall quality of the answers. The chain-based reasoning structure methods (CoA, RAT) achieve win rates ranging from 9.6% to 25.2%, with the win rate for comprehensiveness assessment consistently lower than that of other dimensions. This indicates that while chain-based reasoning structures have a positive effect on answer generation, their lack of multi-branch reasoning capability results in a noticeable disadvantage in comprehensiveness compared to tree-based structures (e.g., RATT) and matrix-based structures (e.g., MTQA). The win rates for the tree-based reasoning structure method (RATT) range from 12% to 34.4%. We attribute this advantage to the column-cell communication mechanism employed by MTQA, which better stimulates the LLMs’ multi-branch reasoning strategies for answering the questions, thereby yielding higher-quality answers.
Ablation Study
Subsequently, we conducted an ablation study and developed four variants of MTQA to demonstrate the effectiveness of each mechanism within MTQA. The comparison is performed using the single-hop QA dataset NQ and the multi-hop QA dataset HotpotQA. The experimental results are shown in Table 3.
No RAG correction: The -RAG variant removes the RAG correction module from MTQA. It is evident that the model’s performance is significantly impacted by the knowledge boundaries and LLM hallucinations, resulting in a notable decline in performance on both datasets. Interestingly, compared to Table 1, the -RAG variant outperforms non-RAG methods, which demonstrates the superiority of the thought matrix over other reasoning structures.
No origin text enhancement: The -origin variant removes the use of the original text within the knowledge unit. It is evident that there is a certain degree of performance loss compared to MTQA. In the multi-hop dataset, the performance of the -origin variant is significantly lower than that of the -KG variant, highlighting the clear advantage of knowledge graph (KG)-structured data in multi-hop problems compared to single-hop datasets.
No KG enhancement: The -KG variant removes the use of knowledge graph triples within the knowledge unit. By combining the data from the -origin variant, it is evident that the absence of certain forms of knowledge leads to a significant performance loss. Furthermore, when handling multi-hop problems, the original text, used as corrective knowledge, has limited effectiveness. We attribute this to the lack of the retrieval capability for structured multi-hop information from the knowledge graph.
No communication: The -comm variant sets the column-cell communication weight matrix in the MoT module to 0, effectively degenerating the MoT structure into a tree-based reasoning structure. From the results in Table 3, it is evident that removing the column-cell communication mechanism causes a significant performance loss for MTQA, even exceeding the performance losses seen in the -KG and -origin variants. This highlights that, during reasoning, the thought structure plays a dominant role compared to the enhanced forms of knowledge.
Dataset Metric NQ HotpotQA
F1 EM F1 EM
-RAG 0.467 0.318 0.384 0.278
-origin 0.489 0.334 0.423 0.294
-KG 0.483 0.329 0.407 0.273
-comm 0.474 0.327 0.409 0.275
MTQA 0.510 0.368 0.452 0.318
Table 3: Comparison of F1 and EM scores of MTQA and its four variant models on two datasets.
Numerical Analysis
In this subsection, we investigate two crucial hyperparameters in the MTQA framework: the column-cell communication weight matrix and the size of the thought matrix. We explore their impact on both the performance and efficiency of the QA task using the HotpotQA dataset.
Weight Matrix Analysis
First, we discuss the impact of the column-cell communication weight matrix settings on the framework’s performance. In this experiment, we conduct a comparative study of 12 different weight matrix configuration methods using a $3\times 4$ matrix as an example. The configuration methods include Uniform, Gaussian, four different constant matrices (0.2, 0.5, 0.8, 1.0), as well as vertically increasing matrices (Vert-0.1, Vert-0.2, incrementing row by row along the y-axis), horizontally increasing matrices (Hor-0.1, Hor-0.2, incrementing column by column along the x-axis), and combined vertical and horizontal increasing matrices (Vert&Hor-0.1, Vert&Hor-0.2, incrementing both row by row and column by column along the x- and y-axes). The Gaussian distribution is set with a mean of 0 and a variance of 1, and the generated matrix is normalized to ensure each weight is between 0 and 1. The evaluation metrics for the experiment are the F1 score and Exact Match (EM) score.
Figure 3: Comparison of F1 and EM scores for various weight setting methods.
The experimental results are shown in Figure 3. We observe that the Vert&Hor-0.1 weight setting method achieves the best performance among all methods, with F1 and EM scores of 0.452 and 0.310, respectively. However, the performance of Vert&Hor-0.2 decreases instead. We infer the following reason: In the early stages of reasoning, the model tends to passively adopt the thought process from the previous thought node, resulting in strategies that are highly similar to the previous ones. Therefore, it is necessary to suppress the influence of the previous thought node’s reasoning in the early stages to further stimulate the LLMs’ reasoning abilities and explore different strategies. The performance of the constant matrix with a weight of 0.2 outperforms that of matrices with weights of 0.5, 0.8, and 1.0, further supporting this viewpoint. In the middle and later stages of reasoning, MTQA has already conducted a deep exploration of the solution space for the problem. However, it may forget the reasoning process and conclusions from the earlier thought nodes, potentially generating strategies that are repetitive. Therefore, it is necessary to expose more conclusions and thought processes, allowing the LLM to actively avoid generating strategies that have already been produced.
MoT Size Analysis
We next discuss the impact of MoT size on its performance and computational overhead. Specifically, we examine how different matrix sizes affect MTQA’s reasoning performance (F1 score and EM) and its time cost (Time). The considered matrix sizes include six configurations: $2\times 2$, $2\times 3$, $3\times 3$, $3\times 4$, $4\times 4$, and $4\times 5$. The weight matrix for each configuration is set using the Vert&Hor-0.1 method described above.
Figure 4: Comparison of F1 and EM scores for different matrix sizes.
As shown in Figure 4, the experimental results indicate that as the matrix size increases, the model’s F1 and EM scores exhibit a gradual upward trend. Specifically, the F1 score increases from 0.392 with a $2\times 2$ matrix to 0.452 with a $3\times 4$ matrix. However, larger matrices ($4\times 4$, $4\times 5$) stabilize around 0.45 or even show a slight decline. The EM score follows a similar trend. This suggests that, for typical multi-hop problems, smaller matrices may lack sufficient depth and breadth in exploring solution strategies; thus, enlarging the thought matrix is an effective way to improve reasoning performance. Nevertheless, this approach is not absolute—once the matrix reaches a certain size, the strategies supported by the available knowledge may have been fully explored. Forcing the LLM to continue exploration at this point may lead to hallucinations and inaccurate answers.
Furthermore, as the matrix size increases, the reasoning time also rises significantly—from 1.324 minutes for a $2\times 2$ matrix to 5.483 minutes for a $4\times 5$ matrix. Therefore, considering both performance and computational cost, the $3\times 4$ matrix achieves the best F1 and EM scores while keeping the reasoning time relatively short, making it a favorable choice. For extremely complex problems or those involving a large volume of knowledge, a larger thought matrix may be considered.
Figure 5: Comparison of MTQA performance under different matrix sizes and different weight setting methods.
Joint Analysis
After obtaining preliminary conclusions for the weight matrix and MoT size individually, we further investigate the interaction between these two hyperparameters within the same experimental framework. The experimental results, shown in Figure 5, lead us to the following conclusions:
Second, in terms of time overhead, regardless of the weight configuration, the reasoning time increases almost monotonically with the number of rows and columns in the matrix. At the $2\times 2$ size, the time ranges from 1.0 to 1.4 minutes, while at $4\times 5$ it rises to approximately 5.4–6.0 minutes. The growth is nearly linear in the early stages, but once the size exceeds $3\times 4$, the time cost increases sharply. In contrast, the F1 and EM scores show no notable improvement, and some weight configurations even result in declines. Therefore, a $3\times 4$ size is the optimal choice in most scenarios.
Overall, the $3\times 4$ configuration combined with Vert&Hor-0.1 is the optimal trade-off: on HotpotQA, it achieves an F1 score of 0.452 and an EM score of 0.318 with an average reasoning time of only 3.2 minutes. This experiment further confirms two key conclusions: (i) the careful design of the weight matrix is more beneficial than blindly enlarging the thought matrix, with small-step vertical–horizontal increments effectively activating diverse reasoning within a column without excessively leaking historical information; and (ii) the thought matrix size has a “sweet spot”—a sufficient number of rows and columns ensures adequate breadth and depth in strategy exploration, but beyond $3\times 4$ the marginal gains diminish while time costs rise sharply. Based on these findings, we adopt $3\times 4$ + Vert&Hor-0.1 as the default MTQA configuration in all comparative experiments above.
Time Complexity Analysis
We compare the time complexity of MTQA with that of the best-performing baseline, RATT, from two perspectives. First, we compare the reasoning time costs of MTQA and RATT on the UltraDomain-Biology and HotpotQA datasets. Second, we provide a theoretical proof of their time costs under the same matrix size configuration.
Method Dataset MTQA RATT
Biology HotpotOA Biology HotpotOA
125 times 402.94 394.56 2804.25 2743.58
average 3.22 3.16 22.43 21.95
Table 4: Comparison of time costs between MTQA and RATT.
As shown in Table 4, both methods were tested 125 times on the Biology dataset. The average reasoning time for MTQA was 3.22 minutes, while RATT required 22.43 minutes. Similarly, on the HotpotQA dataset, the average reasoning time for MTQA was 3.16 minutes, while RATT required 21.95 minutes. The reasoning time of MTQA is approximately one-seventh of that of RATT. This demonstrates that the MTQA method significantly reduces computational overhead and provides efficient reasoning capabilities when handling complex question-answering tasks.
Next, we continue to analyze the time complexity composition of the two methods. Let the constructed MTQA have $m$ rows and $n$ columns, and correspondingly, the RATT tree has $n$ layers, each with $m$ nodes. Let $k$ be the average time to call the LLM to get a response, and $r$ be the time to call RAG. Then:
$T_{RATT}=O(n(2m+1)k)+O(mnr)=O(n((2m+1)k+mr))$ (12)
$T_{MTQA}=O(mnk)+O(2nr)=O(n(mk+2r))$ (13)
In RATT, each layer involves calling the LLM $m$ times to generate thought nodes, followed by $m$ calls to the LLM and $m$ calls to RAG to refine the thought nodes. Finally, a single LLM call is made to generate the summary node. In contrast, in MTQA, we simplify this process by calling RAG once to construct the knowledge unit, followed by $m$ LLM calls to generate the thought nodes enhanced by the knowledge unit. The process is then completed by one call to RAG and one call to LLM for refinement and summarization. Both experimental and theoretical results demonstrate the high efficiency and performance of our approach.
Conclusion
In this paper, we propose the Thinking Matrix Question Answering framework (MTQA) for complex question-answering tasks. The framework, centered around the novel MoT reasoning paradigm, integrates knowledge graph triples and raw text into knowledge units embedded within a retrieval-augmented process. The column-cell communication mechanism triggers multi-branch thinking laterally and accumulates deep strategies vertically, overcoming the common limitations of the CoT’s single path and ToT’s in-layer redundancy. Through extensive experiments across four mainstream datasets, evaluated with multiple metrics and dimensions, we demonstrate the advantages of MTQA over existing methods. Additionally, we reduce the reasoning time to one-seventh of the best baseline, balancing both accuracy and efficiency. This advantage is particularly evident in high-level contextual understanding and other open, abstract tasks.
Limitations: Although MTQA and the core MoT paradigm perform excellently on complex question-answering tasks, there are still several limitations that require further exploration. These include the fact that the effectiveness of the reasoning process and the final responses still depend on the quality of task prompts and the external knowledge base, as well as the high computational demands and the need for manually configured hyperparameters. For future work, we plan to extend the MoT paradigm to various LLM reasoning and alignment tasks, incorporating multimodal information and adaptive communication weight matrices. We aim to validate its generalizability and robustness in broader application scenarios and explore more suitable RAG systems for reasoning structures to achieve higher-quality reasoning enhancement and knowledge correction.
Appendix A Main prompt template
In this section, we present various LLM data processing prompt templates mentioned in the Methodology section, including Figures 6, 7, 9, and 10.
Prompt for graph construct
Figure 6: Graph Construct Prompt.
As shown in Figure 6, the graph construction prompt template is used to build a knowledge graph from text. It consists of several steps: first identifying entities and their descriptions, then extracting relationships between entities, followed by capturing high-level thematic keywords, and finally outputting everything in a structured list. It emphasizes strict formatting without additional explanations or summaries.
Prompt for extract key
Figure 7: Extract Key Prompt.
The prompt template, as shown in Figure 7, is used to extract both high-level and low-level keywords from the user’s query. High-level keywords focus on overarching concepts or themes, while low-level keywords focus on specific entities, details, or concrete terms. The template requires returning the results in JSON format, containing two keys: “high_level_keywords” for overarching concepts or themes, and “low_level_keywords” for specific entities or details. It emphasizes returning only JSON results, without any additional text.
Prompt for initial thought node
Figure 8: Initial Thought Node.
The Initial Thought Node prompt template, as shown in Figure 8, is used to generate structured answers based on previously provided information and current instructions. The template requires the answer to be split into multiple paragraphs and directly respond to the given instructions, without adding extra explanations or introductions. It emphasizes using the previous context and communication information to ensure the structure and consistency of the answer. The output of the template will include context and communication content, following the direct response to the instructions.
Prompt for Summary Node
Figure 9: Initial Thought Node.
The Summary Node prompt template, as shown in Figure 8, is used to generate more detailed and logically coherent answers based on the original answer and an additional supplementary answer. The template requires maintaining a structured format with multiple paragraphs. First, review the original answer, and if any part is correct and does not need further details, retain it unchanged. For parts that need improvement, enhance them by incorporating information from the supplementary answer to make the response more comprehensive and accurate. If any errors are found, they should be corrected while ensuring that the revised content remains logically coherent. Finally, the generated answer should use language consistent with the context and avoid summarizing or replacing words on your own.
Figure 10: example.